find : L inverse [1 divided by (s+1)(s
1
Expert's answer
2021-07-16T03:30:14-0400
"L^{-1}(\dfrac{1}{(s+1)s})=L(\dfrac{1}{s}-\dfrac{1}{s+1})=1-e^{-t}"